SpaceRef

SpaceRef


NASA ISS On-Orbit Status 16 July 2013

Status Report From: NASA HQ
Posted: Tuesday, July 16, 2013

image New Status: USOS Extravehicular Activity (EVA) #23: Flight Engineer (FE)-3 Cassidy (EV-1) and FE-5 Parmitano (EV-2) completed a 1 hour 32 minute EVA with FE-6 Nyberg IVA support. The crew successfully performed the Z1-YJumper install Part 2 and the FGB PDGF 1553 cable final connections. The MLM Ethernet cable was secured to a Node 1 handrail. Due to a buildup of water in Parmitano's helmet, the EVA was terminated early. The crew returned safely to the Airlock and following a nominal repress they egressed their suits and conducted an inspection. EVA specialists are assessing the cause and source of water in the suit.

Water Processor Assembly (WPA) Status: Engineering teams met yesterday to discuss the WPA delta pressure issue. They determined that the next most likely cause is a clogged Mostly Liquid Separator (MLS) Check Valve. The valve is relatively easy to remove but if there is air in the system from the Particulate Filter R&R last Friday there is a risk that the MLS could over pressurize. If there is air in the system, it should be completely dissolved within a week based on current WPA operations. Teams will meet again next week to discuss removing and replacing the check valve.

Mobile Servicing System (MSS) Operations: Robotics Ground Controllers powered up the MSS and maneuvered the Space Station Remote Manipulator System (SSRMS) into position for the EVA task to remove and replace the Lens Cover on the SSRMS End-B Camera Light Pan Tilt Unit (PTU) Assembly (CLPA). It was anticipated that changing the Lens Covers would help determine if the haziness of the image from the CLPA is being caused by the Lens Cover, the CLPA itself or a combination of the two. Following termination of the EVA and Cassidy's and Parmitano's re-entry of the ISS, Robotics Ground Controllers maneuvered the SSRMS back to a park position. MSS performance was nominal.

Today's Planned Tasks

Daily Planning Conference (S-band) [Complete]
Airlock Camera Activation [Complete]
Powerdown the CUCU for US EVA 23 [Complete]
ISS HAM Radio Powerdown [Complete]
USOS Window Shutter Close [Complete]
Comm configuration for EVA [Complete]
Test EVA cameras and transfer to EVA crew [Complete]
Atmosphere Control and Supply (ACS) High Pressure Oxygen Bypass Configuration [Complete]
URAGAN. Observation and Photography using Photo Spectrum System [Complete]
EVA In-Suit Light Exercise (ISLE) Preparation [Complete]
MRM1 Fan Screen Cleaning (group C) СОТР [Complete]
ISS RS ППС-26 and ППС-31 Plug-in Audit [Complete]
ВКС Laptops Antivirus Scan Check and Status Report [Complete]
Life On The Station Photo and Video [Complete]
EMU Purge [Complete]
Prebreathe in EMU [Complete]
Filling (separation) of ЕДВ (КОВ) for Electron [Complete]
Prebreathe in EMU [Complete]
СОЖ Maintenance [Complete]
Crew Lock Depress [Complete]
Post Depress and A/L Egress [Complete]
Crew Lock Egress [Complete]
Connect Cable 1553 [Complete]
Y-Bypass Jumper Installation on Z1 - Part 2 [Complete]
MLM ETHERNET Cable Routing [Partially complete]
V-Guide Bolt Operation on RBG (on P1) [Aborted]
SPDA DOOR [Aborted]
Greetings Video [Complete]
Cleaning LIV TVS Voltage Converter Ventilation Grille [Complete]
WETA Relocation to CP11 [Aborted]
Cleaning ventilation grilles on FGB interior panels (panels 201, 301, 401) [Complete]
Cleaning FGB panel vent screens (116, 316, 231, 431) [Complete]
LCS. Activate RSE-LCS laptop to start data recording [Complete]
MBSU MLI Removal [Aborted]
JEM EF (Japanese Experiment Module - Exposed Facility) Camera Remove and Replace [Aborted]
EVA Get Ahead Tasks [Aborted]
IMS Update Crew
Crew Lock Ingress [Complete]
RWS Powerdown in Cupola [Complete]
Crew Lock Pre-Repress [Complete]
Crew Lock Repress [Complete]
USOS Post-EVA Cleanup [Complete]
Return to nominal comm configuration [Complete]
Reaction Self Test [Complete]
Post-EVA Photo Camera Reconfig [Complete]
Daily Planning Conference (S-band) [Complete]
A/L Camera Deactivation [Complete]
Task List

WHC KTO Replace
WHC EDV R&R
Roscosmos Report
Uragan Report
EKON Photo
Lab PCS HD Swap [In work]
IFM LAB1D5 UAM Check
Regen WSTA Fill
Ground

Nominal Commanding
Three-Day Look Ahead:
Wednesday, 7/17: EMU Airlock deconfig, EVA tool stow, OBT CMO prof training, REGEN ECLSS Recycle Tank R&R
Thursday, 7/18: EMU loop scrub, JSL 4.0 transition, ITCS sample gather/stow,SLAMMD
Friday, 7/19: 34S Emergency Drill, OGS H2 Sensor R&R/AAA Cleaning, EVA Tool Stow

QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up





// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.